OpenStudy (anonymous):

3x^2-3x+4=0 I have to find the discriminant: which is -39 but then I have to determine the type of solutions of the quadratic equation, then solve using the quadratic equation.

jimthompson5910 (jim_thompson5910):

here's a table that may be handy http://www.mathwords.com/d/discriminant_quadratic.htm

jimthompson5910 (jim_thompson5910):

in this case, D = -39

OpenStudy (anonymous):

Yes, I did get that. I mentioned that in my question.... When I put it in the quadratic equation, is 3 going to be negative or positive since I have to plug a -3 into a -b?

OpenStudy (anonymous):

I ended up with x=+ or - the square root of -39 all over 6, but then I get confused

OpenStudy (anonymous):

?

OpenStudy (aum):

\[ x = \frac{-b \pm \sqrt{D}}{2a} = \frac{-(-3) \pm \sqrt{-39}}{2*3} = \frac{3 \pm \sqrt{-39}}{6} = \frac 12 \pm \frac{\sqrt{39}i}{6} \]

jimthompson5910 (jim_thompson5910):

as the table and aum point out, you'll get a complex solution since D = -39 which is less than 0.

jimthompson5910 (jim_thompson5910):

2 complex solutions I mean

OpenStudy (aum):

\[i = \sqrt{-1}\]

OpenStudy (anonymous):

I understand everything except for the last part...why is i suddenly added in

jimthompson5910 (jim_thompson5910):

You cannot take the square root of a negative number. This is because x^2 is never negative. To get around the fact that x^2 = -1 has no real solutions, the idea that \[i = \sqrt{-1}\] was introduced. This is known as an imaginary number. Using this notation, x^2 = -1 now has a solution which is a complex number.

jimthompson5910 (jim_thompson5910):

the solutions to x^2 = -1 are x = i or x = -i

OpenStudy (anonymous):

Thank you very much for your help. I highly appreciate the tutoring.